OpenStudy (anonymous):

MEDAL AND FAN The square pyramid pictured below has a surface area of _____. 144 m 2 120 m 2 189 m 2 126 m 2

OpenStudy (anonymous):



OpenStudy (anonymous):

@hartnn

OpenStudy (mathstudent55):

There are 4 triangular faces and a square base. How do you find the area of a square?

OpenStudy (anonymous):

base x heigh

OpenStudy (aivantettet26):

base x height? i thought it was s^2

OpenStudy (mathstudent55):

A square has 4 sides of equal length. What is the area of a square? You can say length times width, but the length and the width of a square are equal. So to simplify call the length and width the side, or just simply s. \(A_{square} = s^2 \)

OpenStudy (mathstudent55):

In your case, the side of the square base is 6 m. What is the area of the base?

OpenStudy (anonymous):

36 is the area of the square

OpenStudy (mathstudent55):

Great. The area of the square base is 6m^2. Now we find the area of the sides.

OpenStudy (mathstudent55):

Each side is a triangle with a base of 6 m and a height of 9 m. What is the formula of the area of a triangle?

OpenStudy (mathstudent55):

@iribeiro1602 Are you still there?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

so I need to multiply 6 x 9?

OpenStudy (mathstudent55):

np No. What is the formula for the are of a triangle?

OpenStudy (anonymous):

b x h/ 2

OpenStudy (mathstudent55):

You need to calculate the area of the 4 triangular faces. Since all triangles are congruent, you just need to find the area of one of them and multiply it by 4.

OpenStudy (mathstudent55):

Correct. The area of 1 triangle is A = (b x h )/2 = (6 m x 9 m)/2 = ?

OpenStudy (anonymous):

ok so it will be 6 x9 / 2 = 27 x 4

OpenStudy (anonymous):

108?

OpenStudy (anonymous):

108 + 36

OpenStudy (anonymous):

144

OpenStudy (mathstudent55):

Right. 1 triangle has an area of 27 m^2 The 4 triangles have an area of 4 x 27 m^2 = 108 m^2. Now you add that area to the area of the square base, 36 m^2. You got the total correct. Good job!

OpenStudy (mathstudent55):

Btw, nice figure you drew up there!